OpenStudy (anonymous):

2x^2-5x-3? help please. im comfused.. could you factor it?

OpenStudy (cwrw238):

(2x )(x ) the 2 numbers you need in each blank space are obtained as follows they will have a product of -3, and 2 times one of them added to the other will give the middle number -5

OpenStudy (anonymous):

you lost me, could you explain that differently?

OpenStudy (cwrw238):

ok - i'll try - its a bit tricky to explain lets expand a similar expression (2x + 1(x - 2)...................(1) first you multiply second bracket by 2x then by + 1: = 2x^2 - 4x + 1x - 2 = 2x^2 - 3x - 2* now work backwards to get (1) as we have 2x^2 we write (2x )(x ) now to find 2 numbers to go in the blanks: +1 and -2 have product -2* and + 1 + 2(-2) = -3 * so we have what we require to fill the blank spaces (2x + 1)( x - 2)

OpenStudy (cwrw238):